OpenStudy (rayman5766):

Express your answer in lowest terms. 3/8 + 5/16

OpenStudy (anonymous):

Get same denominator 6/16 + 5/16=11/16

OpenStudy (rayman5766):

thanks I have like 25 I did 19 just need help with 2 more

OpenStudy (anonymous):

kk

OpenStudy (rayman5766):

sorry ok here is the question 6/5 - 14/15

OpenStudy (anonymous):

18/15 - 14/15=2/15